Juliana14:

1. If y varies directly as x and y = 2 when x = 3, what is the constant variation?

Juliana14:

Plssss

Juliana14:

With solution

jhonyy9:

ky =x

jhonyy9:

1,5 *2 = 3

jhonyy9:

yeah sure

Juliana14:

What your name

jhonyy9:

ky = x k - is the constant

Juliana14:

What you name on facebokk

jhonyy9:

ky = x k*2 = 3 k = 3/2 1,5*2 = 3 for x = 6 1,5*y = 6 y = 6/1,5 = 4 do you understand it now ?